OpenStudy (anonymous):

the area of a circle is 380 sq inches find the radius

OpenStudy (anonymous):

What is the equation which relates circular area to the radius?

OpenStudy (anonymous):

area = pi X radius squared

OpenStudy (anonymous):

ok, so. \[A = \pi r^2\] \[A = 380 \] \[\pi r^2 = 380\] \[r^2 = \frac{380}{\pi}\] \[r = ?\]

OpenStudy (anonymous):

10.99?

OpenStudy (anonymous):

yeah, thereabouts.

OpenStudy (anonymous):

ohh ok thanks